APPENDIX E: Sample Resume Typeset in Plain TEX

There are a number of “canned” resume style sheets available for use with LaTEX, however it really isn’t too terribly difficult to prepare a a presentable resume using only plain TEX. The following sample illustrates one sort of resume you can readily generate; a machine readable copy of the raw TEX commands used to typeset that resume is available to OREGON VAX users as:

      TEX$DISK:[TEX.SAMPLES]RESUME2.TEX

On other systems get a copy of that file via anonymous FTP.

FTP to DECOY.UOREGON.EDU (128.223.32.19) and see directory pub/tex/samples.

James O. Squeems, III

PineVelCo Timber  226681 Fox Ridge Road
6900 Enterprise Drive  Twin Pines, Oregon 99123
Springfield, Oregon 99234  (503) 555-1234

(503) 555-9876

EMPLOYMENT  Assistant Vice President for Marketing, PineVelCo Timber,

Springfield, Oregon, 1985-date

Senior Marketing Manager, PineVelCo Timber,

Springfield, Oregon, 1979-1984

Marketing Analyst (Veneer Sales), True Timber, Inc., Ely, Minnesota, 1971-1978

Platoon Leader (Infantry), United States Army,

Republic of South Vietnam, 1966-1968

EDUCATION  M. B. A., University of California, Los Angeles, 1970

B. S., University of Minnesota, Minneapolis, 1966

SERVICE  Chairman, Twin Pines Consolidated School Board, 1986-date

Twin Pines Children’s Hospital Board of Directors, 1985-date

AWARDS  Citizen of the Year, Twin Pines, Oregon, 1988

PineVelCo Outstanding Performance Certificate, 1985 and 1980 Certificate of Appreciation, Ely Timbermen, 1976

MEMBERSHIPS  Timber Marketing Association of North America

Twin Pines Country Club

PERSONAL  Married, two children

Active golf and tennis player

Excellent health

REFERENCES  Available upon request